A BACTERIA CULTURE STARTED WITH A COUNT OF 480 AT 8:00 A.M. AND AFTER t HOURS IS EXPECTED TO GROW TO f(t)=480(3/2)t. ESTIMATE THE NUMBER OF BACTERIA IN THE CULTURE AT NOON THE SAME DAYS.
Answer by stanbon(75887)

A BACTERIA CULTURE STARTED WITH A COUNT OF 480 AT 8:00 A.M. AND AFTER t HOURS IS EXPECTED TO GROW TO f(t)=480(3/2)^t. ESTIMATE THE NUMBER OF BACTERIA IN THE CULTURE AT NOON THE SAME DAY.
-------------------
f(t)=480(3/2)^t
From 8:00 AM to noon is 4 hours.
f(4)=480(3/2)^4
f(4)=2430
Cheers,
Stan H.